OpenStudy (aaronandyson):

a body is projected at an angle of 30 to the horizontal with a speed of 30 ms^-1 .The angle made by the velocity vector with the horizontal after 1.5 seconds? g = 10ms-2

OpenStudy (drmoss):

meh

OpenStudy (aaronandyson):

@imqwerty

imqwerty (imqwerty):

Any ideas?

OpenStudy (aaronandyson):

vector law of triangle addition? maine isko resolve kiya

OpenStudy (aaronandyson):

fir us grph ko copy karke x axis ko t leliya

OpenStudy (aaronandyson):

fir 1.5 s pe 0 mila

imqwerty (imqwerty):

wait..grph??

OpenStudy (aaronandyson):

ans ryt h but kya yeh method ryt h?

OpenStudy (aaronandyson):

graph**

imqwerty (imqwerty):

Which graph??

OpenStudy (aaronandyson):

30 ko resolve kiya

imqwerty (imqwerty):

You mean that you resolved velocity into its x nd y components?

OpenStudy (aaronandyson):

yes

imqwerty (imqwerty):

Okay what next

OpenStudy (aaronandyson):

now stuck

imqwerty (imqwerty):

You got stuck here but still managed to get the answer? Anyways assume the velocity in the direction of projection to be v The x component will be vcos(30) which you know is equal to 30 Similarly find y component of velocity Now x component of velocity will remain same cuz no force is acting in x direction but y component will keep changing due to gravity

imqwerty (imqwerty):

Find the velocity in y direction at the given time You already know the velocity in x direction Now you can find the angle between velocity vector and the horizontal by using trigonometry

OpenStudy (drmoss):

what the heck is wrong with your english Aaron xD

OpenStudy (aaronandyson):

@welshfella

OpenStudy (aaronandyson):

the angle is given as 30

OpenStudy (aaronandyson):

how can vcos(30) be same as v?

OpenStudy (aaronandyson):

vcos(30) = 30*sqrt3/2 = 15*sqrt3

OpenStudy (welshfella):

horizontal component of the velocity = 30 cos 30 = 15sqrt3 vertical component = 30 sin 30 = 15 at the point of projection the horizontal component is constant

OpenStudy (welshfella):

at time 1.5 we can can find the vertical component by using v = u - 10t ( im taking acceleration due to gravity to be 10ms-2) v = 15 -10*1.5 = 0

OpenStudy (welshfella):

hmm i did'nt expect that expect that lets check my work

OpenStudy (aaronandyson):

whats wrong?

OpenStudy (welshfella):

a value of 0 means that the angle with the horizontal is 0. Its height is at its maximum

OpenStudy (aaronandyson):

then can the angle be 0?

OpenStudy (welshfella):

of course if we use the more accurate acc due to gravity it will be slightly different ( a = 9.81 ms-2)

OpenStudy (aaronandyson):

Thanks,sir! Can you please help with a few more questions?

OpenStudy (welshfella):

did the question say what value of acc due to gravity to take?

OpenStudy (aaronandyson):

g = 10

OpenStudy (welshfella):

oh ok then 0 degrees is the right answer.

OpenStudy (welshfella):

be back in 5 minutes

OpenStudy (aaronandyson):

Thanks! I have a few more questions i need help with/in.. can you please help more?